Owl Shift summary 99/05/28

Nickolas Solomey (solomey@hep.uchicago.edu)
Fri, 28 May 1999 07:08:11 -0500

Shift Summary for Owl shift, 05/28/99

On Shift: Earl Swallow, Nick Solomey

Others Present from time to time: Jim Graham, Hogan, Arrun,

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of 99 data (special runs not included): 0

Activities, Diversions, and other things like Accounting of lost time:

11:30
NS & ECS arrive. Hall open with many folks working.

NS walks through Exp. Hall -- all is well. He remains in Hall
to work on DC with HN.

01:50
DJ cycled Z-calibrator -- operates OK and left in OUT position.
Vacuum also OK.

JAG found that split writes past end of tape. Will email ST in the
morning.

All of the shift was spent doing pedistal adjustments, the worst
case was the BA and those are mostly done. During the rest of the
time tried to keep the DA running taking cosmics.

No tapes taken to Feynman computer center.

=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=
Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey